OpenStudy (anonymous):

Will give medal and Fan!!(:

OpenStudy (anonymous):

@Chelsea04 @ryan123345

OpenStudy (ryan123345):

okay

OpenStudy (anonymous):

I know that x1 + x2 / 2 , y1 + y2 / 2 is the midpoint formula!

OpenStudy (anonymous):

i just dnt no how to set up

OpenStudy (ryan123345):

(-9+14)/2 (-20+12)/2

OpenStudy (anonymous):

that really means take the average of the coordinates

OpenStudy (anonymous):

The midpoint of a segment uses this formula: M=((x1+x2)/2,(y1+y2)/2)

OpenStudy (anonymous):

the average of \(-9\) and \(14\) is \[\frac{14-9}{2}=\frac{5}{2}\]

OpenStudy (ryan123345):

(5/2,-4) :)

OpenStudy (anonymous):

I like to set my equation up as a coordinate point, but what ryan's saying is right

OpenStudy (ryan123345):

Thank you @Chelsea04 i knew that you couldve handled that problem by yourself :)

OpenStudy (anonymous):

thank you?

OpenStudy (anonymous):

I have a quick question! What it let's say the ending fraction was like, 2/2.

OpenStudy (ryan123345):

that would be 1

OpenStudy (anonymous):

then it'd be 1

OpenStudy (anonymous):

got it! Thanks!!

OpenStudy (anonymous):

always simplify :)

OpenStudy (ryan123345):

No problem. if u need anymore help hmu!